If resistance in three successive arms of balanced bridge is 1, 2 and 36ohms, the resistance of galvanometer in fourth arm will be?

Grade:12th pass

1 Answers

Khimraj
5 years ago
Given that
R1 = 1
R2 = 2
R3 = 36
R4 = ?
and it is balanced bridge then
R1*R3 = R2*R4
1*36 = 2*R4
So R4 = 18 ohm.
Hope it clears. If you like answer then please approve it.